# Tidal heating in binary inspiral of strange quark stars

### v1

@@ V Discussion
We have considered the Newtonian tidal interaction of the binary components as linear perturbations of a background solution for a star. We have analysed the tidal energy dissipated in this bulk-viscous medium from the fundamental flimit-$f-$mode of the star taking into account its dominant coupling with the external tidal field. Since viscosity coefficients are dependent on the stellar temperature, the estimate of the energy dissipation is correlated to the accurate estimation of the temperature during the inspiral. Thermal evolution of the system during the inspiral has been determined considering the dissipated energy and the cooling of the system due to neutrino emission. Regarding the latter, we have computed the cooling timescale employing the neutrino emissivity in unpaired strange quark matter Schäfer and Schwenzer (2004), this being much longer than the inspiral timescale, so it will not have any impact on the temperature estimates.
@@
In order to model unpaired strange quark matter, we have resorted to the non-ideal bag model. We have considered six sets of EoSs with varying strange quark mass $m_{s}$ and the bag constant $B_{\rm eff}$ that span the current EoS uncertainty range and satisfy the astrophysical constraints of maximum observed mass and tidal deformability from GW170817. We have also calculated the heat capacity of quark matter for the EoSs based on the non-ideal bag model to accurately determine the temperature reached at the end of the inspiral. For all the EoSs and equal mass binary systems (of $1.4M_{\odot},1.6M_{\odot}$ and $1.8M_{\odot}$) considered, we have found temperatures of few times $10^{9}$ K reached at the end of the inspiral. These estimates are of the same order as those obtained from hyperonic bulk viscous dissipation (Ghosh et al., 2024) and a few orders of magnitude higher than shear viscosity from $nn$ or $ee$ scattering (Lai, 1994) and bulk viscosity from Urca reactions (Arras and Weinberg, 2019). However, the temperatures are not high enough to require the inclusion of thermal effects in the EoS during the inspiral.
@@
Further, we have used this viscous dissipated energy to estimate the additional phase contribution to the frequency domain gravitational waveform using the single phase approximation. For all the equal mass binary systems considered, the additional phase obtained is $\sim 0.05-0.1$ rad and $\sim 0.1-0.5$ rad for sets where $m_{s}=150$ and $200$ MeV, respectively. Since it is expected that nonperturbative effects at high densities in QCD can contribute to make the strange quark mass higher, further constraints on this parameter and their impact on the bulk viscosity calculation are a line of improvement in the estimation of the phase shift in the gravitational waveform.
@@
We have found that this extra phase shift in the gravitational waveform is higher than the expected upper limits of the phase uncertainty for nearby sources (at distances of $100$ Mpc) considering current state-of-the-art waveform models and the proposed future generation GW detector sensitivities. Therefore, this additional phase should in principle be detectable with next generation detectors for high SNR events. It should be noted that the inferred phase uncertainties only serve as approximate upper limits of this effect and should be incorporated into waveform models to obtain accurate estimates. When compared with the estimates for hyperonic bulk viscous dissipation, we get a detectable phase difference for $1.4M_{\odot}$ and $1.6M_{\odot}$ binaries of strange quark stars. For hyperons, this happens only for very massive systems ($\geq 1.8M_{\odot}$), as hyperons appear at high densities achievable only at the core of massive stars. Therefore, the signature of tidal dissipation during binary inspiral in low mass binary systems ($\sim 1.4M_{\odot}$) can be a ‘smoking gun’ signature for the strange quark stars.
@@
There are several future directions that can be explored based on this work. Firstly, more consolidated efforts should be given to develop state-of-the-art gravitational waveform models that incorporate the dissipative aspects of tides. Recent works have incorporated this dissipative effect in frequency domain models (Ghosh et al., 2025) or in terms of ‘tidal lag’ in effective theory (Ripley et al., 2023, 2024; Saketh et al., 2024), but these studies should be improved to accurately capture the tidal dissipation over the entire parameter space of mass and frequency during inspiral. Secondly, it would be interesting to determine the amount of dissipation in hybrid stars as well, although we expect this to be much smaller than strange stars given the smaller quark core. Finally, it is also instructive to consider other phases of strange quark matter, whether in the core of a hybrid star or as strange quark star. For example, the bulk viscosity has already been studied for color-superconducting strange quark matter in the 2SC phase Alford and Schmitt (2006); Alford et al. (2024), spin-one color-superconducting strange quark matter Wang and Shovkovy (2010), and color-flavor-locked phases Alford et al. (2007); Manuel and Llanes-Estrada (2007); Alford et al. (2008b); Bierkandt and Manuel (2011); Mannarelli and Manuel (2010), being of particular interest the 2SC phase, which produce a similar bulk-viscous profile in the low-temperature regime as the normal phase Schmitt and Shternin (2018).
@@
To conclude, tidal dissipation effects during binary inspiral offers a unique possibility to probe transport properties of dense matter where conservative tidal effects (adiabatic and dynamical) only probe the equilibrium EoS of dense matter. For low mass compact binary systems ($\sim 1.4M_{\odot}$), a detection of tidal dissipation in gravitational waves can be a ‘smoking gun’ signature for the presence of strange quark stars since there is no known nuclear matter source of high dissipation in that regime. Given their astrophysical significance and considering the increased sensitivity of next generation GW detectors, accurate BNS waveforms should be developed in near future incorporating tidal dissipation effects.
@@ Acknowledgments
The authors thank Swarnim Shirke, Nils Andersson and Alexander Haber for helpful discussions. SG acknowledges support from STFC via grant no. ST/Y00082X/1. We also acknowledge support from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), as well as from the Generalitat de Catalunya under contract 2021 SGR 171, and by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211. B.K.P acknowledges the use of IUCAA Pegasus cluster for the computational work and also the support of the CNRS-IN2P3 MAC masterproject, and the project RELANSE ANR-23-CE31-0027-01 of the French National Research Agency (ANR). This work makes use of NumPy van der Walt et al. (2011), SciPy Virtanen et al. (2020), Matplotlib Hunter (2007), jupyter Kluyver et al. (2016), and Corner software packages.
@@ Appendix A Solving for the $f$-mode frequency within Relativistic Cowling Approximation
Non-radial oscillation modes have been studied for many decades. The procedure for analysing these modes in the non-relativistic framework was first discussed by Cowling, while in the context of general relativity, Thorne and Campollataro explored this further. In general relativity, it is necessary to include the metric perturbation to solve the perturbed fluid equations. However, within the Cowling approximation, the metric perturbations can be neglected. It was demonstrated that the oscillation frequencies of the $f$-mode obtained using the Cowling approximation and those calculated using the complete linearized equations of general relativity differ by less than 20%. As outlined in the main manuscript, we will adopt the Cowling approximation, where the spacetime metric for a spherically symmetric background is given by:
$ds^{2}=e^{2\Phi(r)}dt^{2}+e^{2\Lambda(r)}dr^{2}+r^{2}d\theta^{2}+r^{2}\sin{\theta}^{2}d\phi^{2}.$ (39)
@@
$\frac{dW(r)}{dr}$ $=$ $\frac{d\epsilon}{dp}\left[\omega^{2}r^{2}e^{\Lambda(r)-2\phi(r)V(r)+\frac{d\Phi(r)}{dr}W(r)}\right]$
@@
$\frac{dV(r)}{dr}$ $=$ $2\frac{d\Phi(r)}{dr}V(r)-\frac{1}{r^{2}}e^{\Lambda(r)}W(r),$ (40)
where,
$\frac{d\Phi(r)}{dr}=-\frac{1}{\epsilon(r)+p(r)}\frac{dp}{dr}.$ (41)
The functions $V(r)$ and $W(r)$ along with frequency $\omega$, characterize the Lagrange displacement vector $\mathcal{\zeta}^{i}$ associate to a perturbed fluid
$\zeta^{i}=\left(\frac{W(r)}{e^{\Lambda(r)}},-V(r)\partial_{\theta},-\frac{V(r)}{\sin^{2}{\theta}}\partial_{\phi}\right)\frac{Y_{lm}(\theta,\phi)}{r^{2}},$ (42)
where $Y_{lm}(\theta,\phi)$ is the $lm$-spherical harmonic. A solution of Eq. (40) with the fixed background metric in Eq. (39) near the origin will behave as follows:
@@
$\omega^{2}e^{\Lambda(R)-2\Phi(R)}V(R)+\frac{1}{R^{2}}\frac{d\Phi(r)}{dr}\Big{|}_{r=R}W(R)=0.$ (44)
@@ Appendix B Neutrino emissivity of quark matter
According to the discussion in Ref. Schäfer and Schwenzer (2004) and the work done in Ref. Heiselberg and Pethick (1993) about the thermal conductivity, it can be assumed that degenerate quark matter is isothermal because the equilibration is fast. As a result, the cooling taking into account only neutrino emissivity and neglecting surface emission is given by
$\frac{dU}{dt}=C_{V}(T)\frac{dT}{dt}=-\mathbf{\epsilon}(T),$ (47)
then
$t-t_{0}=-\int_{T_{0}}^{T}\frac{C_{V}(T)}{\mathbf{\epsilon}(T)},$ (48)
where we assume that the chemical potentials’ variations with the temperature are negligible (see the appendix in Ref. Hernández et al. (2024)). Thus,
$T$ $=$ $\left(Ct+T_{0}^{-4}\right)^{-1/4}.$ (49)
with
$C=\frac{914}{105}\frac{G_{F}^{2}\cos^{2}{\Theta_{C}}\alpha_{s}\mu_{u}\mu_{d}\mu_{e}}{N_{c}\left(\mu_{u}^{2}+\mu_{d}^{2}+\mu_{s}k_{s}\right)+\mu_{e}^{2}},$ (50)
where light quarks and electrons are considered to be massless.
In Fig. 6 we show the thermal evolution of strange quark matter employing Eqs. (49),(50). We determine the chemical potentials at $n_{B}=2.5n_{0}$ using the EoS from perturbative QCD up to $\mathcal{O}(\alpha_{s})$ and setting the renormalization scale at $\bar{\Lambda}=2\mu_{s}$ Fraga and Romatschke (2005). Considering massless light quarks and the running of strong coupling and the strange quark mass we obtain $\alpha_{s}\approx 0.7$ and $m_{s}=161$ MeV. As can be seen, the cooling due to neutrino emission is negligible for temperatures below $2.5\cdot 10^{9}$ K in a timescale of $1000$ s. Therefore, its contribution is small for the study of the tidal heating of strange quark stars.

### v2

@@ VI Discussion
We have considered the Newtonian tidal interaction of the binary components as linear perturbations of a background solution for a star. We have analysed the tidal energy dissipated in this bulk-viscous medium from the fundamental $f$-mode of the star taking into account its dominant coupling with the external tidal field. Since viscosity coefficients are dependent on the stellar temperature, the estimate of the energy dissipation is correlated to the accurate estimation of the temperature during the inspiral. Thermal evolution of the system during the inspiral has been determined considering the dissipated energy and the cooling of the system due to neutrino emission. Regarding the latter, we have computed the cooling timescale employing the neutrino emissivity in unpaired strange quark matter Schäfer and Schwenzer (2004), this being much longer than the inspiral timescale, so it will not have any impact on the temperature estimates.
@@
In order to model unpaired strange quark matter, we have resorted to the non-ideal bag model. We have considered six sets of EoSs with varying strange quark mass $m_{s}$ and the bag constant $B_{\rm eff}$ that span the current EoS uncertainty range and satisfy the astrophysical constraints of maximum observed mass and tidal deformability from GW170817. We have also calculated the average heat capacity of quark matter for the EoSs based on the non-ideal bag model to accurately determine the temperature reached at the end of the inspiral. For all the EoSs and equal mass binary systems (of $1.4M_{\odot},1.6M_{\odot}$ and $1.8M_{\odot}$) considered, we have found temperatures of few times $10^{9}$ K reached at the end of the inspiral. These estimates are of the same order as those obtained from hyperonic bulk viscous dissipation (Ghosh et al., 2024) and a few orders of magnitude higher than shear viscosity from $nn$ or $ee$ scattering (Lai, 1994) and bulk viscosity from Urca reactions (Arras and Weinberg, 2019). However, the temperatures are not high enough to require the inclusion of thermal effects in the EoS during the inspiral.
@@
Further, we have used this viscous dissipated energy to estimate the additional phase contribution to the frequency domain gravitational waveform using the single phase approximation. For all the equal mass binary systems considered, the additional phase obtained is $\sim 0.06-0.2$ rad and $\sim 0.1-0.5$ rad for the sets where $m_{s}=150$ and $200$ MeV, respectively. Since it is expected that nonperturbative effects at high densities in QCD can contribute to make the strange quark mass higher, further constraints on this parameter and their impact on the bulk viscosity calculation are a line of improvement in the estimation of the phase shift in the gravitational waveform.
@@
We have found that this extra phase shift in the gravitational waveform is higher than the expected upper limits of the phase uncertainty for nearby sources (at distances of $100$ Mpc) considering current state-of-the-art waveform models and the proposed future generation GW detector sensitivities. Therefore, this additional phase should in principle be detectable with next generation detectors for high SNR events. It should be noted that the inferred phase uncertainties only serve as approximate upper limits of this effect and should be incorporated into waveform models to obtain accurate estimates. When compared with the estimates for hyperonic bulk viscous dissipation, we get a detectable phase difference for $1.4M_{\odot}$ and $1.6M_{\odot}$ binaries of strange quark stars. For hyperons, this happens only for very massive systems ($\geq 1.8M_{\odot}$), as the size of hyperon core increases with the mass of neutron stars. Therefore, the signature of tidal dissipation during binary inspiral in low mass binary systems ($\sim 1.4M_{\odot}$) can be a ‘smoking gun’ signature for the strange quark stars.
@@
There are several future directions that can be explored based on this work. Firstly, more consolidated efforts should be given to develop state-of-the-art gravitational waveform models that incorporate the dissipative aspects of tides. Recent works have incorporated this dissipative effect in frequency domain models (Ghosh et al., 2025) or in terms of ‘tidal lag’ in effective theory (Ripley et al., 2023, 2024; Saketh et al., 2024), but these studies should be improved to accurately capture the tidal dissipation over the entire parameter space of mass and frequency during inspiral. Secondly, it would be interesting to determine the amount of dissipation in hybrid stars as well, although we expect this to be much smaller than strange stars given the smaller quark core. Finally, it is also instructive to consider other phases of strange quark matter, whether in the core of a hybrid star or as strange quark star. For example, the bulk viscosity has already been studied for color-superconducting strange quark matter in the 2SC phase Alford and Schmitt (2006); Alford et al. (2024), spin-one color-superconducting strange quark matter Wang and Shovkovy (2010), and color-flavor-locked phases Alford et al. (2007); Manuel and Llanes-Estrada (2007); Alford et al. (2008b); Bierkandt and Manuel (2011); Mannarelli and Manuel (2010), being of particular interest the 2SC phase, which produce a similar bulk-viscous profile in the low-temperature regime as the normal phase (see Fig. 5 in Ref. Schmitt and Shternin (2018)).
@@
To conclude, tidal dissipation effects during binary inspiral offer a unique possibility to probe transport properties of dense matter where conservative tidal effects (adiabatic and dynamical) only probe the equilibrium EoS of dense matter. For low mass compact binary systems ($\sim 1.4M_{\odot}$), a detection of tidal dissipation in gravitational waves can be a ‘smoking gun’ signature for the presence of strange quark stars since there is no known nuclear matter source of high dissipation in that regime. Given their astrophysical significance and considering the increased sensitivity of next generation GW detectors, accurate BNS waveforms should be developed in near future incorporating tidal dissipation effects.
@@ Acknowledgments
D.C. would like to thank Prof. J. Schaffner-Bielich for pointing to her the article by Hernandez et al (2024) which led to this work. The authors thank Swarnim Shirke, Nils Andersson and Alexander Haber for helpful discussions. SG acknowledges support from STFC via grant no. ST/Y00082X/1. We also acknowledge support from the program Unidad de Excelencia María de Maeztu CEX2020-001058-M, from the project PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER, UE (FSE+), as well as from the Generalitat de Catalunya under contract 2021 SGR 171, and by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211. B.K.P acknowledges the use of IUCAA Pegasus cluster for the computational work and also the support of the CNRS-IN2P3 MAC masterproject, and the project RELANSE ANR-23-CE31-0027-01 of the French National Research Agency (ANR). This work makes use of NumPy van der Walt et al. (2011), SciPy Virtanen et al. (2020), Matplotlib Hunter (2007), jupyter Kluyver et al. (2016), and Corner software packages.
@@ Appendix A Solving for the $f$-mode frequency within Relativistic Cowling Approximation
Non-radial oscillation modes have been studied for many decades. The procedure for analysing these modes in the non-relativistic framework was first discussed by Cowling Cowling (1942), while in the context of general relativity, Thorne and Campollataro Thorne and Campolattaro (1967) explored this further. In general relativity, it is necessary to include the metric perturbation to solve the perturbed fluid equations. However, within the Cowling approximation, the metric perturbations can be neglected. It was demonstrated that the oscillation frequencies of the $f$-mode obtained using the Cowling approximation and those calculated using the complete linearized equations of general relativity differ by less than 20%. As outlined in the main manuscript, we will adopt the Cowling approximation, where the spacetime metric for a spherically symmetric background is given by:
[…]
@@
$\frac{dW(r)}{dr}$ $=$ $\frac{d\epsilon}{dp}\left[\omega^{2}r^{2}e^{\Lambda(r)-2\nu(r)}V(r)+\frac{d\nu(r)}{dr}W(r)\right]$
@@
$\frac{dV(r)}{dr}$ $=$ $2\frac{d\nu(r)}{dr}V(r)-\frac{1}{r^{2}}e^{\Lambda(r)}W(r),$ (47)
where
$\frac{d\nu(r)}{dr}=-\frac{1}{\rho(r)+p(r)}\frac{dp}{dr},$ (48)
with $p$ being the pressure and $\rho$ the energy density as stated above.
The functions $V(r)$ and $W(r)$ along with frequency $\omega$, characterize the Lagrange displacement vector $\mathcal{\xi}^{i}$ associated to a perturbed fluid
$\xi^{i}=\left(\frac{W(r)}{e^{\Lambda(r)}},-V(r)\partial_{\theta},-\frac{V(r)}{\sin^{2}{\theta}}\partial_{\phi}\right)\frac{Y_{lm}(\theta,\phi)}{r^{2}},$ (49)
where $Y_{lm}(\theta,\phi)$ is the $lm$-spherical harmonic.
A solution of Eq. (47) with the fixed background metric in Eq. (46) near the origin will behave as follows:
@@
$\omega^{2}e^{\Lambda(R)-2\nu(R)}V(R)+\frac{1}{R^{2}}\frac{d\nu(r)}{dr}\Big|_{r=R}W(R)=0.$ (51)
@@ Appendix B Neutrino emissivity of quark matter
The cooling in a strange quark star due to neutrino emissivity and neglecting surface emission can be studied assuming the medium to be isothermal given that thermal equilibration in degenerate matter is fast Schäfer and Schwenzer (2004); Heiselberg and Pethick (1993); Schmitt and Shternin (2018). Thus, the thermal evolution equation is given by
$C_{V}\frac{d\tilde{T}}{dt}=-L,$ (54)
where the average neutrino luminosity $L$ is computed over the entire stellar volume Ofengeim et al. (2017)
$L(\tilde{T})=\int_{0}^{R}\epsilon(\rho,\tilde{T})e^{2\nu}\frac{4\pi r^{2}dr}{\sqrt{1-2m/r}}.$ (55)
In the isothermal approximation, the average neutrino luminosity can be written as the following equation
$L=\upsilon\tilde{T}^{6},$ (56)
where $\upsilon$ depends on stellar volume and the equal-mass configuration. Using the simplified equations for the average heat capacity (see Eq.(32)) and neutrino luminosity in the isothermal approximation, we obtain an expression for the final redshifted temperature, $\tilde{T}_{f}$
$\frac{1}{\tilde{T}_{f}^{4}}=\frac{1}{\tilde{T}_{0}^{4}}+\frac{4\upsilon t}{\sigma}.$ (57)
Here $t$ is the time that takes the system to reach a final temperature $\tilde{T}_{f}$ from an initial temperature $\tilde{T}_{0}$.
@@
Figure 6: Thermal evolution of strange stars. The final redshifted temperature is normalized to $10^{9}$K and is shown as a function of the time (in seconds) for five initial temperatures ($\tilde{T}_{0}=$$5\cdot 10^{8},\,10^{9},\,2.5\cdot 10^{9},\,5\cdot 10^{9},\,10^{10}$ K), with $\alpha_{s}\approx 0.31$ and $m_{s}=150$ MeV.